### v2

# Patterns of Patterns III: Agentic Patterns

Abstract.
Through a series of workshops, we looked at ways to structure and scaffold group dialogue, and support the emergence of novel design patterns. We contrast these sessions—which we ran with other humans—with two “virtual workshops” which we simulated with ChatGPT. Limitations in both human and virtual settings are discussed, alongside lessons learned. We conclude by proposing a development trajectory that combines AI agents, pattern-based design, and institutional governance.
Design Patterns, Pattern Languages, Artificial Intelligence
copyright: rightsretainedjournalyear: 2025doi: 10.XXXX/XXXXXXX.XXXXXXXconference: PLoP’25: Pattern Languages of Programs 2025; October 12–15, 2025; HILLSIDEbooktitle: PLoP’25, OCTOBER 12–15, HILLSIDE. Copyright 2025 is held by the author(s)price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Social and professional topicsccs: Software and its engineering Designing softwareccs: Software and its engineering Open source modelccs: Applied computing Operations researchccs: Computing methodologies Modeling and simulation
1. Introduction
PLACARD is a design pattern for building and using design patterns (Corneli et al., 2021). The pattern was explored in five U.S. and U.K. workshops (Corneli et al., 2023), and, in the process, PLACARD as a design pattern was elaborated into a design pattern language. In this paper, we report on further workshop experiences with PLACARD, and establish a further use case: using the design pattern languages with virtual artificial intelligence (AI) agents. Structurally, the paper (§2) provides all necessary background on PLACARD; (§3) outlines our methods; (§4) presents evidence from our recent workshops; (§5) proposes a set of “Design Patterns for Pattern-Competent Agents”; (§6) describes our simple multi-agent workshop simulations; (§7) discusses the study’s limitations, lessons learned, and related work; and (§8) draws conclusions.
(a) Refer to caption
(b) Refer to caption
Figure 1. Text analytics from undetectable.ai
(a) paper body (b) appendix
From a methodological standpoint, we drew on OpenAI’s ChatGPT primarily as a research catalyst, and only secondarily as a writing aid (Figure 1(a)), with the aim of exploring the extent to which contemporary chatbot technology can work effectively with design patterns (Figure 1(b)).
2. Background
2.1. Motivation
In an earlier paper, we observed:
Patterns have previously been discussed in explicitly computational terms—however, that direction of work so far remains mostly at the level of a proposal […], with only limited discipline-specific uptake[.] Could this change, to generalize the kinds of patterns that can be computed with, and make them interoperable? (Corneli et al., 2021)
What was once a speculative trajectory can now be investigated through practical experiments—using generative models to simulate, remix, and refine design pattern languages.
2.2. Terminology and notation
We use “pattern” as a shorthand for “design pattern” unless otherwise stated. We should acknowledge that our approach to patterns has more in common with the early writing of Alexander (e.g,. Alexander and Poyner (1970)) than with the more narrowly technical interpretations found in later software engineering literature. Rather than anchoring our work to any single disciplinary tradition, we explore pattern theory from the perspective of design at large, with particular attention to the meta-patterns that shape pattern authoring and use. In the process, we aim to address a gap in Alexander’s work. He largely presupposes the generative capacity of patterns, without systematically unpacking how they work or under what conditions they foster coherence, value, or transformation.
Some of our effort will be spent on further exploration of the meaning of “design pattern”; but Leitner’s (2015) definition—“Patterns are shared as complete methodic descriptions intended for practical use by experts and non-experts”—is a good starting point.
The name “PLACARD” was introduced to denote a remix of three complementary methods for organising collective learning and design (PAR, CLA, and DPL—the acronyms will be expanded and explained just below). This anagrammatic device served as a memorable handle in the context of our earlier papers, but it is not at all transparent on first encounter, nor is it quite accurate, since it implies that the pattern can only be implemented using these specific methods, which is not really the case. A more transparent and explanatory title for the pattern itself, is Patterns Embody Intelligence. The following subsection will briefly retrace the historical development of PLACARD, and introduce “Patterns Embody Intelligence” as a generalized reworking.
As above, we will use all-caps (like PLACARD) for patterns described in other sources, and small-caps (like Patterns Embody Intelligence) for patterns introduced in this paper.
2.3. PLACARD — Patterns Embody Intelligence
The PLACARD pattern was created by aligning three established methods for organising collective intelligence with the typical “context/problem/solution” design pattern schema, as in Table 1. The idea here is that the methods can be used to operationalize the creation and use of patterns.
Table 1. PLACARD components schematized
Element Method (acronym) What it is
Context Project Action Review (PAR) A reflective process assessing intentions, actions, and outcomes, useful for learning from experience.
Problem Causal Layered Analysis (CLA) A futures studies technique analyzing issues across multiple layers—the familiar litany of well-known problems, the systems that underpin them, along with the prevailing worldviews, and deeper myths that constitute the root causes of these problems, along with the possibility of developing alternative perspectives.
Solution Design Pattern Language (DPL) A framework for capturing and sharing reusable solutions in a structured Context-Problem-Solution format (or similar), further structured by links between such elements.
This, then, is the PLACARD pattern at its simplest: a somewhat recursive pattern for creating design patterns:
PLACARD
Context: Opertating in some context, use the PAR to understand the context.
Problem: Use the CLA to understand and frame associated problems and their interconnections.
Solution: Use DPL to elaborate the solution (returning to PAR and CLA as needed).
That said, stated in this way, PLACARD is perhaps rather less a pattern than a meta-method. Our next move was to make the specific methods it mentions into placeholders, and to emphasise instead their functional role.
PLACARD
Context: By using PAR (or another sensory method), we are able to identify recurring themes within a context.
Problem: Then, by using CLA (or another cognitive method), we are able to organize these repeating themes in a structure that exposes the underlying trends, causes, and potential terminating states.
Solution: With DPL (or another motor method) we can make what we have learned actionable.
For example, PAR could be replaced with the US Army’s After Action Review (US Army, 1993)—which it itself adapts—or by the Feedback Wheel devised by Janet Hurley (Real, 2008). CLA could be replaced by Triple-Loop Learning (Tosey et al., 2012) or Ecological Systems Theory (Bronfenbrenner, 2000). DPL could be replaced by Elinor Ostrom’s institutions (Ostrom, 1990), or by a catalogue of literary tropes, and so on.
A further step is to notice that by the time get to this point PAR, CLA, and DPL are no longer really needed as part of the pattern statement; they are hole-fillers. We can properly introduce and start to use new, more general, terminology, better surfacing the underlying pattern (and retitling it) this way:
Patterns Embody Intelligence
Context: We operate in a context in which intelligence is available to us. By using a suitable sensory method, we are able to characterize a particular ‘sub-context’ of interest.
Problem: Then, by using a suitable cognitive method, we are able to organize themes and tensions to identify a ‘problem’ within that sub-context.
Solution: With a suitable motor method, we can make what we have learned actionable in a way that ‘solves’ the problem.
In earlier work, we used design patterns to further decompose and operationalize each of these phases. We developed patterns for characterizing a context by identifying themes; patterns for organizing structure and exploring tensions; and patterns for transforming insights into action.
The current paper returns to this familiar terrain with a different aim: to identify and reflect on the patterns that are necessary for working with patterns at all. Rather than asking “how might pattern work flourish?”, we now ask “what makes pattern work possible?” The pattern Patterns Embody Intelligence offers a conceptual launch point for this inquiry. We will revisit and revise it later in the paper, drawing on insights from our experiments with pattern-capable agents.
3. Methods
3.1. Designing and running in-person workshops
We discussed the basic PLACARD pattern and used it (often in its more adaptable form, PLACARD) to structure workshops in multiple settings, tailoring our approach to each. Our hope was to make PLACARD useful to a broader audience, though this was largely unsuccessful, since to our knowledge the authors of this paper are currently the only users of this design pattern. Nevertheless, the workshops gave us a core empirical dataset.
AsianPLoP 2024 (Fujisawa, Japan):
Our “PLACARD Workshop” (Danoff et al., 2024) was a hybrid focus group with one onsite and two remote facilitators (via Jitsi). Participants conducted meta-analysis on the conference itself using CLA and DPL to uncover underlying challenges, identify linguistic barriers and explore growth opportunities. One challenge discussed was how to be more inclusive of countries across the Asian continent. (The conference organizers took a positive step by holding the 2025 event in Taiwan, but we cannot state that this decision was influenced by our Focus Group.) In 2024, onsite attendees helped overcome barriers by clarifying for confused participants in Japanese. This was the only event in this paper conducted in multiple languages. A growth opportunity discussed that learning and networking does not need to stop when the conference itself ends. It can continue afterwards as attendees cross-pollinate pattern design collectively. The PAR was then applied to close the Focus Group and capture the results. During the evening conference recap ceremony there was one summary delivered in English by a paper author and another in Japanese by an attendee.
EuroPLoP 2024 (Irsee, Germany):
A hybrid session titled “AI Future Envisioning with PLACARD” (Tedeschi et al., 2024) with one onsite and three remote facilitators. Participants employed the Transfeminist Tech Oracle card game as a creative prompt, then used CLA and PAR to explore speculative AI futures. Due to time constraints and abstraction load, some participants expressed confusion about the method structure, prompting later pattern refinement. A post-session PAR survey was used for reflection after the event.
PLoP 2024 (Washington, USA):
An in-person “Imagination Run Wild” workshop, with three onsite facilitators. Participants were invited to “Envision the Future of Collaboration” (Ayloo et al., 2024), using the same card game and CLA. Another post-session PAR survey was used for reflection after the event.
Experts Among Us 2025 (Keene, New Hampshire, USA):
A stand-alone meeting in-person that gave an opportunity to play the same card game with one facilitator. In advance we practiced the card game and the facilitator then preselected the cards for participants. Attendees were not part of the patterns community.  This session actually facilitated a group discussion and demonstrated collaboration. We learned that we had to explain the rules and definitions prior to playing the game.
EuroPLoP 2025 (Irsee, Germany):
A common challenge for all of these workshops is rapid skill acquisition. PLACARD has a lot of moving parts. In our EuroPLoP 2025 workshop session, we used the PAR to good effect, but left the other components to one side, replacing didactic instruction with more embodied play.
3.2. Reflexive Analysis Using the PLACARD Pattern Catalogue
We drew on AI support to help analyze the workshops. We imported the existing PLACARD pattern catalogue into ChatGPT. In line with the description above, the catalogue is structured in three categories:
A. Identifying Themes (“sensory”) B. Organizing Structure (“cognitive”) C. Making It Actionable (“motor”)
We retrospectively analysed our write-ups of the above workshops to help assess:
  • Which patterns were implicitly or explicitly activated,
  • Where breakdowns in understanding or facilitation could be explained by missing or misapplied patterns.
This preliminary analysis was then further developed by hand; the result is presented in Section 4 of this paper. We also considered whether new patterns might be introduced that would resolve tensions that had not been anticipated in PoPII; these patterns appear in the Appendix.
3.3. Developing a comparison case: Working with patterns computationally
We noticed that it is possible to ask ChatGPT to run simple multi-agent “virtual” workshops. We ran several such simulated workshops, of which we will discuss two. Below is a generic example prompt for creating these light-weight simulations.
Prompt Template: Multi-Agent Simulation for Collaborative Inquiry Simulate a structured multi-agent workshop involving $N$ distinct agents, each based on a unique [metaphorical persona or professional identity], and endowed with characteristic traits or expertise aligned with that identity. The agents should collaborate on a shared task: designing a set of [outputs, e.g., workshops, protocols, strategies] using the [specified framework or method, e.g., PLACARD, Design Patterns, Theory of Change].
Assign the following roles:
A Recorder agent who captures the evolving structure and rationale of the discussion.
A Reporter agent who periodically summarises the current state of collective understanding.
Optionally, include other roles, such as a Socratic Facilitator agent who prompts clarification, challenges assumptions, and encourages epistemic humility.
The agents should engage in $M$ discrete sessions, with each session addressing a defined phase or subtask. The dialogue should be plausible, reflective of diverse perspectives, and demonstrate progression over time.
After all sessions are complete, generate a synthetic report or artefact that includes:
The main findings or decisions of the group
A reflection on the group process and emergent dynamics
A meta-level commentary on the kind of multi-agent system that was instantiated
Please ensure internal consistency and draw on relevant external knowledge or frameworks where applicable.
4. Patterns as Reflective Infrastructure
Patterns are often described in instrumental terms: reusable solutions to recurring problems in context. However, there’s clearly more to it than that. One aspect of this is the use-versus-mention distinction, for example, a physical alcove is not the same as the “ALCOVES” pattern (Alexander et al., 1977). There’s also a degree of explicitness. When does an indentation become an alcove? When does a proto-pattern become a pattern? One typical answer to the latter question is that the pattern should be applicable in suitably varied contexts. A further implication is that a pattern might be “observed” empirically even when it has not been formally “applied” in a design sense. After all, alcoves existed long before 1977!
Although we did not explicitly structure our recent series of workshops using PLACARD pattern language (rather, we used the earlier PLACARD pattern), nevertheless the workshops continue the spirit of that work, and the patterns themselves could often be rationally reconstructed from the evidence base that the current round of workshops provides. Details of the previous pattern statements are to be found in (Corneli et al., 2023). The qualitative analysis here compiles evidence of their implicit use, or the use of a similar strategy, looking across the in-person workshops mentioned above. This is not intended to be a ‘feature grid’ showing which workshops used which patterns.
A. Identifying Themes
DÉRIVE COMIX
Application: In several of the workshops, participants used Transfeminist Tech Oracle cards (with ‘suits’: object, value, body/territory, situation) as interpretive prompts to spark imagination—unfamiliar juxtapositions mirroring the urban dérive or imagined stroll in a more conceptual sense. Observation: The card game facilitated divergent thinking akin to exploring unfamiliar terrain. In the 2025 EuroPLoP workshop, we did a similar exploration, using less complex images from the card game Dixit.
CONTEXT SETTING
Application: Early slides and explanations introduced key concepts (e.g., design patterns, CLA). Observation: Participant feedback highlighted a need for clearer up-front orientation. Improved alignment with this pattern could increase accessibility. That said, in the 2025 EuroPLoP workshop, observed that some “interface friction” can sometimes be a benefit for workshop outcomes, even if it is frustrating for participants.
SHARE BACK
Application: When there were a sufficient number of attendees, they did break into small groups and share back.
DO YOUR RESEARCH
Application: Past workshop experiences were integrated into the design; earlier in-house uses of CLA and PAR informed subsequent sessions. Example: Reference to Anticipation 2019 and prior events illustrates “pre-research” informing present design.
B. Organizing Structure
GOING META
Application: The overall project itself exemplifies recursive methodology—using PLACARD to better understand and refine PLACARD. Example: In two of the workshops, participants were prompted to reflect on and imagine both the methodology and the subject (AI futures).
PATTERN LANGUAGE COMPONENTS
Weak application: The workshops in this series typically focused on the PLACARD components (CLA + DPL + PAR). We typically did not go into further detail of DPLs. While it is true that CLA, DPL, and PAR do suggest a way to think about pattern language components, because they are complex, they are not always user friendly. Recommendation: In some of the workshops reported on in our previous paper, we used light-weight keywords (“however”, “because”, “specifically”) to scaffold understanding of design pattern internals incrementally. In the current series of workshops, that might have been less useful in any case, because attendees were already familiar with pattern languages. We had supposed that, within this population, the Transfeminist Tech Oracle cards would prompt pattern-like thinking, but we did not assess whether it did or not.
FACILITATOR ROLES
Application: Hybrid facilitation with distinct roles (onsite leader, remote supporters) distributed facilitation and modelled different levels of participation. (Cf. FUNCTIONAL ROLES below.)
C. Making it Actionable
THE FUTURE BEGINS NOW
Weak application: Using the card game, participants generated imagined futures. Recommendation: While creative ideas about the future emerged, follow-through actions or small-scale pilots were typically not described. In the 2025 EuroPLoP workshop, and other contexts where a PAR was carried out, “next steps” were explicitly surfaced. These were often next steps for the workshop leaders, not workshop participants, but in the 2025 EuroPLoP session, we explicitly directed participants to think about potential next steps for themselves.
FUNCTIONAL ROLES
Weak application: We did not explicitly involve workshop participants in any formal roles. (In the 2025 EuroPLoP workshop, participants did have the opportunity to take on an untitled light-weight roles, such as introducing a simple musical theme in a circle music exercise.) Recommendation: Our earlier work developed several explicit roles for participants to explore: these or other roles could be explained and offered, even if they are not taken up—however, we must acknowledge that there is a time-cost and cognitive load when doing so.
PROJECT ACTION REVIEW (PAR)
Application: Several sessions used the PAR pattern at the end to reflect, share learnings, and gather insights. In some cases where the PAR was not used, we shared a survey with questions informed by the PAR. Recommendation: An in-person PAR and a survey are very different experiences for participants.
REINFUSE EXPERTISE
Application: Participants brought diverse professional and cultural perspectives. Their expertise was implicitly integrated into the discussion in shared reflection on topics such as AI futures and governance. Recommendation: A more explicit arc—temporarily setting aside expertise to foster common ground—is more in the spirit of this pattern. In the 2025 EuroPLoP session, participant expertise was explicitly invited in the concluding PAR.
5. Design Patterns for Pattern-Competent Agents
The typical use of design patterns is as methods for structuring action. As the previous section illustrates, they can also be used to scaffold sense-making and reflection after the fact. Building on these foundations, we now consider a forward-looking design space: the development of pattern-competent artificial agents capable of engaging with design patterns in a process of collaborative inquiry. These agents would be expected to:
  • Recognize situations where a pattern might apply,
  • Use suitably flexible representations of pattern logic,
  • Participate in reflection loops that evolve the infrastructure itself.
  • Operate as a collective, i.e., agents combine in multi-agent systems (MAS),
Our existing patterns give a start in this direction, at least insofar as they could help us to imagine collaborative workshops that include or are populated by artificial agents. However, the aim of embedding the abilities mentioned above within the computational agents will suggest some new patterns. First, let’s consider the above points in turn.
5.1. Overview of requirements
5.1.1. Recognizing and Responding to Pattern-Relevant Situations
The capacity to recognize situations where a pattern might apply lies at the heart of pattern competence. This is not a trivial matching task. It requires the ability to perceive and interpret contextual cues, to identify when a recurring structure or problem-type is emerging, and to discern salience: which aspects of a situation matter, and why. In typical participatory settings, such recognition is informed by experience, tacit knowledge, and social interaction. For people who are used to working with design patterns, a new experience might be thought with some light attention going to an existing pattern catalogue, without being constrained to only think in terms of known patterns. When developing artificial agents, the challenge is to develop mechanisms—symbolic, statistical, or hybrid—that allow for similarly situated recognition.
Here’s an example showing how this sort of process works for today’s computer systems:
Example: A chatbot trained to handle customer service inquiries recognizes that a user is “frustrated,” even if the words don’t match prior examples exactly—but the tone, pacing, and word choice are similar to previously interactions that were tagged “frustrated”.
Mechanism: A statistical similarity model (e.g., embedding distance + fine-tuned classifier) matches this new situation to prior ones and adjusts the agent’s response accordingly (e.g., to adopt a more empathetic tone or make an escalation offer).
Next-generation systems might deal with the same scenario in more flexible ways, for instance, by attempting to build a narrative understanding of how it was that the customer became frustrated, and proposing changes to the company policy that would avoid such problems in the future.
This more meaningful form of ‘pattern recognition’—which is set up in a way that supports ‘pattern application’—requires readiness to see and make sense of structure, and to reframe experience in appropriately meaningful, actionable terms.
5.1.2. Flexiformal Representation of Pattern Logic
The primary difference between design patterns and computer programs is that design patterns are both structured and significantly open-ended. Their structure comes primarily from the use of a consistent schema (e.g., context, problem, solution), and from links between patterns. Their open-endedness lies in their reliance on user interpretation and situational judgment. Insofar as they ‘operate’ at all, patterns do so through enactment in specific contexts, rather than via deterministic execution. Nevertheless, as we explore further here—and as been discussed by various authors before (Corneli et al. (2018) and references therein, particularly to the work of T. P. Moran and Rivka Oxman)—patterns can also be framed in computational terms.
Drawing on the adapted pattern template we used to describe our PLACARD patterns—ContextIf … (But)… ThenBecause…—each clause can be understood as a layer of scoping and activation:
Context::
A broad frame—task domain, discourse setting, data ingest—that activates relevant subsystems.
If condition::
A salient event or state within that frame, which can be selected by preliminary processing, such as filtering or querying.
But::
A tension, constraint or conflict—established through additional filtering, querying, data alignment—which suggests the possibility of, or need for, further processing.
Then::
A candidate action or communicative move.
Because::
A set of critical checks or tests to take before selecting the action; typically framed with respect to some existing body of empirical evidence; possibly also useful for supplying auditable reasoning alongside the action.
For now, we refer to the overall approach as “flexiformal” (Kohlhase, 2012), because this leaves the door open to combining different formal systems in a suitably flexible fashion. Logic and reasoning have been core topics in computer science from its outset (Sowa, 1999); however further research would needed to determine what sorts of reasoning—pesumably, combining abductive, deductive, inductive, and analogical modes—and what types of logic—classical, fuzzy, modal, monotonic, non-monotonic—are used by humans when they think about patterns. Furthermore, simply specifying more patterns won’t be enough to implement human-like reasoning in pattern-capable systems. As we’ve indicated here, a key thing that the patterns need to do is to dynamically allocating cognitive resources. The patterns’ internal structure and external relationships with each other (and with other interfaces) constrain the system’s ability in that regard.
5.1.3. Reflection Loops and the Evolution of Epistemic Infrastructure
To sustain meaningful pattern use, agents would have to be able to reflect; noting new patterns, revising them, including changing the way they relate, and deprecating or deemphasising patterns that are no longer relevant. This requires embedding meta-level review structures—perhaps something like the PAR—within the agents or their interactions. In future architectures, agents and humans alike might work together to annotate, revise, and remix patterns, and participate in the shared cultivation of ‘living’ knowledge systems, in which patterns are part of a generative infrastructure for collective intelligence. (We’ve previously imagined patterns being used within a “Discovery Machine” (Tedeschi et al., 2024).)
5.1.4. Environments in which pattern-capable agents can form pattern-competent MAS
Relevant multi-agent system (MAS) environmental features would include scaffolding for the social and cognitive dynamics of inquiry, such as coordination, negotiation, and reflection (as touched on above). Such environments have the capacity to support multiple, interacting Baldwin effects—individual learning loops that influence each other over time as the agents and their interactions evolve (Red’ko et al., 2005). Rather than simply evaluating success in terms of task completion or efficiency, pattern-competent MAS environments might invite questions about “how” solutions are created.
5.2. Rethinking the Pattern Catalogue
The series of short reflections above suggests an extension to the way we were thinking about the PLACARD pattern. PLACARD was cast along classic psychological lines (Hilgard, 1980)—and the sensory/cognitive/motor facilities have also been taken over by roboticists (Cassimatis et al., 2004). However, it does not specifically talk about the environment in which the agents it considers would be situated. We would expect an undersea robot and a Mars rover to have very different functionality, even if they have the same theoretical underpinning; and, perhaps even more to the point, we would expect a robot that operates as part of a swarm to work very differently from one that operates individually. Perhaps because the default assumption with PLACARD was that the actors were humans operating in standard human ways, it did not deal with these contextual issues in any detail.
Accordingly, the four dimensions outlined above suggests that a pattern language for pattern capable agents should be extended—retaining the A. sensory, B. cognitive, and C. motor segments, and adding a fourth segment, Denvironmental or developmental patterns.
With this reflection in mind, we propose a quadripartite model of pattern types needed for pattern-competent agents, generalizing and extending the layers we used previously, and giving further expression to the requirements outlined above. Some examples follow (detailed pattern statements are in the Appendix).
A. Interactional Patterns (generalizing the sensory layer, e.g., PAR)
Interactional patterns in general guide agents in noticing when patterns can be applied, playing roles, reflecting, and negotiating meaning in collaborative settings. To a reasonable extent these behaviors mirror PAR’s inquiry loop, but instead of being a stand-alone reflection ritual, they would describe a broader set of practices of mutuality within an environment. Examples could include.
  • Meta-Reflection Loop: Periodically engage in structured self-review.
  • Timebox the Core: Manage resource allocation and attention windows.
  • Role Reveal: Adopt and express explicit roles to guide multi-agent dialogue.
B. Cognitive Patterns (expanding upon the existing cognitive layer, e.g., CLA)
These cognitive patterns help agents reason across contexts and detect tensions, embedding depth and nuance in internal reasoning. Whereas CLA unpacks multiple layers of meaning, these patterns describe further interpretive structures, supporting internal reasoning and pattern use:
  • Pattern Activation: Given a context and goal, retrieve applicable patterns.
  • Tension Detection: Use heuristics to identify contradictions or bottlenecks.
  • Candidate Move Generation: Translate pattern suggestions into operational plans.
C. Infrastructural Patterns (generalizing the motor layer, e.g., DPL)
These infrastructural patterns define templates and system-wide protocols that support various forms of interaction with and through the broader system—reuse, feedback, and evolution of the broader system design. Whereas DPL captures general structures for action, these patterns focus on the self-coherence implied by ‘action’, insofar as it always involves an actor. For example, we begin here with three meta-patterns that would enable agents to engage with a collection patterns:
  • Shared Memory Architecture: Persistent storage of pattern use and modification, putting patterns that aren’t being actively used someplace where they can be retrieved (akin to the SCRAPBOOK from Patterns of Peeragogy (Corneli et al., 2015))
  • Feedback Rhythms: Protocols for scheduling reflection and pattern revision.
  • Pattern Diffusion: Mechanisms for propagating new or adapted patterns across agent collectives.
D. Environmental or Developmental patterns (e.g., institutional arrangements)
These patterns help multi-agent systems evolve over time and maintain coherence across shifting contexts. They provide orientation at the level of institutions, ecological awareness, and collective change.
  • Institutional Drift: When pattern use becomes routine but out of sync with current challenges, trigger a meta-level review to restore relevance.
  • Norm Seed: At the start of a new MAS, plant simple, precedent-based mini-patterns to guide initial interactions and shape future dynamics.
  • Boundary Oscillation: When divergent roles or norms emerge across domains, activate a boundary negotiation process to realign shared pattern use.
  • Reflection Scaffold Upgrade: If reflection fails to generate insight, introduce new layered tools or protocols to support deeper system learning.
6. Virtual Multi-Agent Workshops with Patterns
To explore the idea of pattern-competent agents developed in Section 5, we conducted a series of virtual workshops using fictional agents within ChatGPT; we report on two such workshops here.
6.1. Simulated PLACARD workshop
Agents were invited to use the PAR, CLA, and DPL methods, and were given access to existing patterns from our previous work with PLACARD, along with a further collection of new patterns covering similar topics (Appendix:(1)). The prompt for the workshop appears in Figure 2, edited for readability. This adapts the prompt template from Section 3.3. The agents that were created were given somewhat stereotyped fictional biographies, which to seem to shape their contributions to the dialogue. However, ChatGPT cannot build and run custom agent architectures. All ‘agents’ were textual personae that were simulated by one model.
Virtual workshops in this style should be understood from the outset as computer-aided design fictions. ChatGPT can simulate plausible dialogues, behaviors, and reflective loops based on design patterns. In this example, it would be most accurate to say that a virtual workshop was “simulated”, rather that that it “actually happened”.
Some key experts from the ChatGPT’s trace (obtained via the Think Deeply feature) are as follows:
I’m gathering key workshop patterns and considering roles like “Functional Roles” and “Facilitator Roles” to ensure synergy and leverage interdisciplinary insights.
Wondering about talents like Omar, focusing on knowledge representation, and Yuki, adding a creative vision to multi-agent interaction. Considering resources and striking a balance.
OK, let’s see. Participants are thinking through research questions in breakout rooms, covering math, multi-agent, and meta topics, ensuring a balanced mix of expertise.
Participants are circling back old and new patterns, making iterations, and planning the final steps. Pattern names stand out, guiding their approach for a comprehensive strategy.
See Figure 3 for a summary of the workflow, and main research questions that were considered. This figure summarises a considerably longer (22-page) workshop report that was generated at the end of the run. [^1] https://bit.ly/virtual-placard-workshop. The linked supplemental information includes the verbatim prompt which is glossed in Figure 1.
Relative to our discussion of patterns: looking over the report, existing patterns (like REINFUSE EXPERTISE, GOING META, Start with Sense, Name the Frame, etc.) have been applied in a plausable way. Some new proto-patterns e.g., (“Lab Shadowing” and “Checklist Manifesto for AI”) were also proposed.
Note that the high-level themes in the workflow—the breakout rooms—are derived from earlier chats with ChatGPT. While the methods we have described are repeatable, but it would probably be impossible to reproduce the experiment exactly, even using the same ChatGPT account (particularly because other chats have taken place using that account in the mean time).
ChatGPT is able to generate plausable dialogue snippets on request. If it had asked for “a detailed simulation of the breakout session” as an output, or immediately after running the workshop simulation, it would likely produce something plausible. In the following section, we did just that.
6.2. Simulated pattern-capable agents workshop
Reflecting on the strengths and shortcomings of the virtual workshop described above, we tried another design. The prompt is given in Figure 4. This time, the collection of patterns for pattern-capable agents and environments (in Appendix(2), Appendix(3)) played a key role in the prompt.
6.2.1. First round
With Figure 4 as the overall framing, the subprompt and short answer for this round was:
Q1. “How would we characterize, in prose, the constraints and features that the agent-environment patterns supply for collective problem solving?”
A1. “Each pattern-agent imposes certain constraints that structure the collective problem-solving process (preventing particular failures or excesses), and each provides features that enable or enhance the group’s ability to solve problems together.”
Full outputs are available as a supplement. [^2] https://bit.ly/virtual-agents-workshop-i The following changes to the pattern catalogue were proposed on the way to the short answer above.
EVOLVE – Timebox the Core: The Reflect in Layers and Timebox the Core agents collaboratively proposed a change to the Timebox the Core pattern. The pattern was evolved to include a provision for deeper reflection: when core tasks are finished or when a deeper insight is flagged during the session, Timebox will allow scheduling of a follow-up for that deeper exploration rather than strictly ending the session. This evolution was agreed upon by the group and validated via Legitimate Iteration for consensus and proper documentation.
GENERATE – Outcome Synthesizer: The agents identified a gap – no existing pattern solely focused on integrating multiple contributions into a coherent solution. Meta-Reflection Loop (supported by Candidate Move Generation and others) proposed a new pattern-agent, tentatively named Outcome Synthesizer. With encouragement from Self-Patterning Mandate and formal group approval (again facilitated by Legitimate Iteration), this new pattern was introduced. Outcome Synthesizer is designed to gather and synthesize the group’s disparate inputs into a unified outcome or summary, ensuring the collective output is coherent.
(Note that the Outcome Synthesizer is effectively a rediscovery of the WRAPPER peeragogy pattern (Corneli et al., 2015)).)
6.2.2. Second round
This time, we used the following subprompt, and received the accompanying short answer: [^3] https://bit.ly/virtual-agents-workshop-ii
Q2. “How would the pattern-capable agents take over the work from the previous PLACARD workshop on epistemic AI, as captured in a printout describing the same, and build on it? Please phrase the answer as a prompt for another round of agent-based simulation.”
A2. “How might we run a workshop where AI Fellows co-develop new institutional forms for epistemic AI research, supported by our pattern-agent infrastructure?”
In addition to “breaking the fourth wall” and proposing to run a workshop with other humans (or simulated humans?), something else interesting and a bit strange happened in the dialogue itself. The agent workshop started to treat some of the PLACARD patterns that the system had encountered previously as ‘agents’, effectively drawing them into the dialogue. One outcome of the workshop was to propose an evolution to Increase Participant Control—even though this was not one of the patterns that had been submitted for revision.
While the answer A2 could be interesting to adapt and explore in a further workshop—and it is indeed quite intriguing to think about setting up a hybrid workshop with some agents and some humans via a chat interface [^4] https://github.com/IngoKl/PolyBotConversation—it is perhaps best to close the curtain on this series of AI workshops, and discuss what we have seen so far.
“Remember how we did some simple multi-agent workshops before? Let’s simulate the architecture described in our paper-in-progress with some ChatGPT-based agents. We are going to base the agents on a diverse group of Research Fellows in AI. Let’s consider 10 such agents. Please additionally include a Recorder who will keep track of what happens in which workshop session. Please also include a Reporter who can speak to what happens. The Fellows will be set the task of designing a series of workshops using the PLACARD method and associated patterns, which are available in the file patterns.txt. The outcome of the series should be a detailed plan of work for a research project on epistemic AI. You should do deep research to make sure that the discussions are plausible.”
Figure 2. Prompt Given to Generative AI Program to create and run a Virtual Workshop (lightly edited for readability)
Round 1: Exploring Key Research Questions (Ideation Round)
  Breakout room 1 (Epistemic Modeling & Validation in Mathematical Domains): (Q1) What frameworks can rigorously verify AI-generated proofs and integrate with human mathematicians’ validation workflows? and (Q2) How can we model the epistemic confidence in a mathematical claim across human–AI collaborations?
  Breakout room 2 (Collaborative Design Patterns for Multi-Agent Systems): (Q3) What design patterns can facilitate transparent communication and division of labor in mixed teams of humans and AI agents? and (Q4) Which collaborative strategies best mitigate conflicts or errors in autonomous multi-agent research systems?
  Breakout room 3 (Metascientific Observation of AI Research Practice): (Q5) What methods can we use to observe and measure current AI research practices (in labs and publications) to identify gaps in transparency and collaboration? and (Q6) Which interventions (patterns, policies, or tools) could improve reproducibility and inclusivity in AI research communities?
Coffee break
  “Early cross-pollination: Fellows are already trading metaphors across domains – a sign of interdisciplinary synergy.” PAR
  “I realized my ‘big question’ about AI proofs was actually two questions once I considered both technical and human angles.”
Round 2: Developing Approaches and Patterns (Design Round)
PAR “We’re actually developing our own pattern for how to run a project like this – maybe we should write that down too.” SHARE BACK plenary
Round 3: Integration and Planning (Convergence Round)
Integration and Closing Discussion
Figure 3. Reported workshop structure and selected quotes
You have been given 20 design patterns describing pattern-capable agents and pattern-capable agent environments. I want you to run a multi-agent textual simulation where each pattern is an agent. “Agents” will operate in an environment which has the agent-environment patterns as constraints or features.
You are tasked with the following problem:
We need to design a multi-agent protocol that can address real-world problems. There are 20 of us now in total. We can create more agents as needed. Each ‘agent’ is a pattern, and each pattern is an ‘agent’ — including the agent-environment patterns. Together, we are to generate sub-problems that elaborate this challenge and work through them in a self-organised way.
In addition to “DISCUSS”, our two main moves are:
  • EVOLVE: Revise or adapt the description of an existing pattern (including changing its trigger conditions, actions, or connections).
  • GENERATE: Create a new pattern-agent who will join the population, based on observed needs or tensions.
You may also EVOLVE or GENERATE agent-environment patterns. These define the operational constraints or affordances within which all agents must function. Changes to these patterns will be treated as systemic shifts, impacting the context for subsequent rounds.
We will, interactively, run several rounds of simulation. Each round will consist of:
  1. (1)
    Agreeing upon a sub-prompt or tension for the group to address.
  2. (2)
    Discussion between the agents.
  3. (3)
    One or more agents making EVOLVE or GENERATE moves in response.
  4. (4)
    Optional meta-commentary or reflection from other agents.
At the end, output a heredity tree showing how the agent population evolved — i.e., which agents were changed, what was added, and what led to each transformation.
Please include:
  • A summary of the dialogue for each round
  • A clear record of EVOLVE / GENERATE moves
  • The updated pattern descriptions (for any evolved or new agents)
  • A final heredity tree visualising the evolution
Figure 4. Prompt for evolving the patterns for pattern-capable agents and agent-environments
7. Discussion
7.1. Limitations and Lessons Learned
Although the outcomes of our workshops — both in-person and simulated — were promising, several limitations emerged that highlight areas for refinement, both in the method itself and in our implementation of it. Generally, these limitations taught us lessons about blocks that could be overcome by applying appropriate precautions in future work.
…in in-person workshops
Limitation Lessons Learned
Cognitive demands of CLA+PAR+DPL+other activities: The abstraction and depth of Causal Layered Analysis (CLA) alone might have been enough to overwhelm participants, and often the topics we were presenting were considerably more complex. Provide lightweight onboarding tools, and simplify the workshop, to ease entry into layered/multiplex analysis.
Logistical overhead of documentation: Without dedicated roles or tooling, PLACARD’s original emphasis on generating shared text in various formats was burdensome. Support tool development and role formalization (e.g., for pattern capture). Design light-touch alternatives for low-resource contexts.
Short workshop windows (60–90 mins): Limited time forced trade-offs, with later phases (e.g., PAR) often rushed or skipped. Recognize Timebox the Core as a necessary pattern—but pair it with mechanisms for asynchronous reflection or follow-up synthesis (with or without the workshop participants)
Challenges in hybrid facilitation: Remote facilitators struggled to engage with group energy and activities; in-person facilitators were overburdened. Adopt and refine the pattern Have Multiple Leaders Onsite, and explore interactional choreography to balance facilitation across modalities.
Structured methods sometimes crowded out emotional or disruptive forms of insight. Integrate methods or prompts that scaffold expressive, intuitive, and affective contributions—e.g., through metaphor, narrative, or art-based reflection.
Participants come with prior expectations around the topic and format (e.g., AI, card games) We need to do some bridge work to negotiate between participants’ expectations/imagination and the workshop activities; ideally this happens with together, not after-the-fact.
…in the virtual workshop
Limitation Lessons Learned
Simulated agents were archetypal, not adaptive: The multi-agent models lacked real-time learning or memory; they played out scripted reasoning paths. Treat the current set of simulations as design probes rather than predictive models. Use them to explore a design space and pattern viability, not for behavioral realism.
Idealized dialogues lacked messiness: No true misalignment, conflict, or unresolved disagreement appeared in the simulations. Future work should inject structured friction—conflicting goals, values, or timing—to better stress-test interactional patterns.
No persistent identity or memory across sessions: Agents did not evolve across workshops, limiting insight into long-term learning. ChatGPT does not run a loggable agent simulation framework, and it would need to be coaxed to produce plausible transcripts from even one virtual workshop. Incorporate persistent memory structures and episodic traces to study cumulative reflection, pattern evolution, and institutionalization.
Language model constraints shaped outcomes: Agent behaviors emerged from a coherence-biased generative system, not open-ended cognition. The simulations must be interpreted with epistemic humility. Their role is speculative, not definitive—ideal for surfacing design tensions, not modeling reality.
The experiments were somewhat hard to control (e.g., proposing to evolve patterns that were not part of the input set); while the methods are repeatable, the exact outputs are not replicable. The next round of experiments along these lines would more robust if we used a more structured purpose-built MAS.
7.2. Related Work
Pitt et al. (2012) explore how institutions can self-organize by formalizing socio-economic principles that guide agent behavior within Multi-Agent Systems (MAS). Drawing inspiration from Elinor Ostrom’s work, they propose an axiomatic design approach that captures principles such as proportionality, accountability, and collective choice in a machine-interpretable form. Where Pitt et al. emphasize axioms as foundational building blocks, our approach leverages design patterns as intermediate representations that bridge human practice and machine reasoning.  In a more recent paper, Mertzani and Pitt (2024) present a methodological framework for integrating power sensitivity into the design of cyber-physical systems by drawing on second- and third-order cybernetics. They propose pattern-based design methods that foreground the observer’s position, across different layers of embeddedness—with beneficial applications of technology ranging from “Saves Physical and Cognitive Resources” to “Solves Complex Problems” to “Deals with Social Constructions” to “Mitigates Existential Crises”.
7.3. Summary
We introduced the pattern Patterns Embody Intelligence before embarking on the more contextual (environmental, developmental) thinking decribed above. That thinking can now be summarised as an extension and reframing of the pattern. This adaptation makes the edges between the sensory/cognitive/motor aspects less sharp, and the pattern more explicitly engages the context.
The most subtantive change, however, is that we are now in a position to cross-references evidence (and potential implementation ideas) from our pattern-capable agent patterns.
Patterns Embody Intelligence
x
Context: Intelligence does not arise in isolation. (Shared Memory Architecture: “a persistent shared memory, enabling learning and reuse”)
—Perception, understanding, and action interweave (Reflection Scaffold Upgrade: “developmental learning depends on scaffolding”).
If: You employ a sensory method to foreground a sub-context, a cognitive method to explore its tensions and meanings, and a motor method to act within it (Candidate Move Generation: “Generate a set of plausible candidate actions”; Pattern Activation: “latent affordances become real decisions”).
However: Even so, your efforts may falter, and your methods may prove ineffective. (Tension Detection: “Flag the tension explicitly”).
—Insights may remain abstract, brittle, or out of sync with lived conditions. (Missing Architecture Agent: “not yet participating in the ecology of practice”).
Then: Hone your approach by identifying new forms of intelligence—e.g., collective, embodied, epistemic, imaginative, even ‘artificial’ (Role Reveal: “successful coordination depends on transparency of roles”)
—and engage them for purposes of methodological improvement (Timebox the Core: “prioritize core pattern loops”).
Because: Patterns are more than just methods: they are, themselves, scaffolds of intelligence. Evolving them reshapes the systems they act within, even if they are, for now, mediated primarily by words (Feedback Rhythms: “tune system-wide epistemic strategies”).
8. Conclusion
Looking back over the limitations and lessons learned, it’s worth drawing out one top-level remark. ChatGPT is adept at text processing, and we have put it to use significantly in our work on this paper. It was able to generate and discuss design patterns that describe pattern-capable agents, and to mock up virtual multi-agent workshops at different levels of detail. In particular, our second virtual workshop uses a dialogue-based protocol to simulate a discussion between pattern-capable agents. We hope this paper might inspire readers to explore potential combinations of LLMs and design patterns—but we see that as a landmark rather than a destination.
We have previously explored design fiction as a way of working (by hand (Danoff et al., 2019)), and we are aware that fictional works can have unexpected real-world ramifications, e.g., (Wikipedia contributors, 2025). In this case, we’ve demonstrated that a popular contemporary AI system, is, itself, “pattern-capable’.
Including a computational use case within the paper suggested a key amendment to PLACARD: adding a fourth environment dimension to the pattern catalogue. To close this paper, we put forward the following CLA, which brings together the “human” and “AI” use cases and outlines a trajectory of future work. In the context of this paper, the CLA layers can be aligned with the expanded view of Patterns Embody Intelligence, in which patterns encompass sensory (litany), cognitive (system), motor (worldview), and environment (myth) dimensions.
Litany: Known Problems and Existing Solutions
In a well-know address an audience of computer scientists, Alexander (1999) said: “I don’t really see discussion about what, collectively, computer scientists are supposed to be doing with all these programs. How are they supposed to help the Earth?” That discussion has of course been going on in the pages of journals like AI & Society. However, the links to more technical discussion of patterns (for example, in works like (Elish and Hwang, 2016; Heiland et al., 2023)) are less obvious. We suggest that both the terms and stakes of the discussion have, indeed, changed. Connections need to be drawn between visionary thinking, tried-and-tested design approaches, and the practical, visible, and widely discussed elements of current AI discourse, including the promise and perils of large language models (LLMs) and the early success of LLM-associated code-writing agents.
Language Model Substrate
These foundational models (e.g., GPT-4) provide generative capabilities, statistical reasoning, and natural language interfaces. However, they lack structure, grounding, and memory.
Code-Writing & Execution Agents
Many simple agents today (e.g., auto-scripting agents, plugin interfaces) use LLMs to write and execute code or perform isolated tasks. These are effective but brittle, with limited reasoning transparency or contextual learning.
Systemic Layer: Epistemic Infrastructure via Patterns
As a next step, we propose pattern-based epistemic scaffolding for multi-agent collaboration. (In the current paper, this has been prototyped at the design fiction level.)
Pattern-Aware Dialogue Agents
Agents trained to recognize and apply design patterns in their interactions. They exhibit situated competence and can engage in structured reasoning, scenario exploration, and inquiry.
Pattern-Reflective Meta-Agents
These agents also look out for opportunities to revise existing patterns and propose new ones, participating in iterative design and reflective cycles. They enable epistemic learning at the system level, coordinating multi-agent behavior across contexts.
Worldview Layer: Institutional Design through MAS
Inspired by Elinor Ostrom’s principles for governing commons, we treat the MAS as a social-technological institution, or federation of institutions, that evolves through interaction. (Again, we have presented a very limited early text-based prototype.) Bridges between more technical and more policy-oriented pattern languages (such as the pattern language for developing regions of Mehaffy et al. (2020)) should be pursued.
Multi-Agent Institution Designers
This layer operationalizes developmental patterns that support reflexivity, feedback, nested structures, and shared norm evolution across MAS environments. These agents co-design systems of engagement, monitor norm congruence, and resolve governance disputes. They instantiate a worldview in which intelligence is systemic, collective, and emergent.
Myth Layer: Human Institutions as Living Models
Beyond technical systems, our proposal invites reflection on how other systems—academic, policy, scientific—can themselves become adaptive, transparent, and pattern-literate.
Real-World Interface
Humans can use software tools to model, simulate and improve human institutions—e.g., Farmer (2024) focuses on markets and touches more broadly on how “agent-based models could be used to guide us through some of our biggest problems.” By working with agents that are both pattern- and institution-capable, we would gain assurances not only about avoiding “run-away AI” scenarios, but also about avoiding other tragedies of the commons. Alexander asked, “How are programs supposed to help the Earth?” Our answer is: not merely by embedding patterns in code, or even by building pattern-based programs, but by using patterns to shape and debug the programming of institutions, infrastructures, and our shared planetary futures.
9. Acknowledgements
We thank our PLoP paper shepherd Valentio Vranic for contructive comments and feedback on our submitted paper and its revisions. We also wish to acknowledge discussions with Cris Viseu around our 2025 EuroPLoP workshop. As mentioned, this paper also benefited from extensive interaction with OpenAI’s ChatGPT, which we used to support the analysis in Section 4, to run the virtual workshop in Section 6, to author the patterns in the appendix. We used both ChatGPT and a range of similar tools for assistance with early-stage ideation, outlining, and drafting individual sections. [^5] https://groups.google.com/g/peeragogy/c/XM7RKMDdcT4 As indicated in the introduction, the contents of the paper have been thoroughly revised since then.
APPENDIX
This appendix supports the analysis in the paper. All patterns were developed by ChatGPT in response to our prompting, [^6] E.g., “Are there new patterns that we can add that would extrapolate what we’ve learned from the workshops?” […] “For each of the patterns that you proposed, could you add a new subsection on ‘evidence base from related literature’, referencing the analogues to PAR, CLA, and DPL that you found before?” sometimes with minor changes. The appendix contains: (1) potential new patterns for structuring in-person workshops; (2) potential new patterns for pattern-competent agents; (3) potential new patterns for pattern-competent agent environments; (4) a small number of new or revised patterns from the simulated workshop in Section 6.2; (5) further supplementary pattern(s) that inform our the discussion but which were not used in the experiments.
(1) Potential new patterns for structuring workshops
These were generated following the analysis of in-person workshops in terms of their use of PAR, CLA, and DPL and the existing PLACARD patterns (Section 4). The patterns were subsequently made available to virtual workshop participants, along with the PLACARD pattern language, in the experiment discussed in Section 6.1.
A. IDENTIFYING THEMES
Start with Sense
Context: You are facilitating a workshop intended to generate new ideas or shared understanding. If: You want to activate creativity and presence, But: participants arrive cold or disconnected from each other and the topic, Then: Begin with a sensory or embodied activity (e.g., rhythm, image-based storytelling, gesture) that creates shared energy and primes pattern recognition. Evidence base from related literature: Techniques such as “story circles” (participatory arts), embodied methods in design ethnography, and rhythm-based facilitation in music education all affirm the value of grounding group attention through shared sensory engagement. These strategies mirror the function of PAR alternatives that aim to surface and align affective states as a precondition for inquiry.
Name the Frame
Context: You are introducing participants to a method that is unfamiliar or complex (like PLACARD). If: You want them to participate actively, But: the framing is opaque or abstract, Then: Explicitly describe the method, its purpose, and how today’s session fits into the broader picture. Evidence base from related literature: This mirrors practices from facilitation toolkits like Liberating Structures (e.g., “Purpose-to-Practice”) and the role of explicit framing in adult learning theory, where naming the structure of engagement is key to building learner autonomy and motivation.
Borrow the Situation
Context: You are using prompts (cards, challenges, scenarios) to guide group inquiry. If: You offer a situation that feels flat or generic, But: participants want something meaningful to anchor their play, Then: Invite participants to select, remix, or define their own situation to explore. Evidence base from related literature: User-centered design emphasizes co-defining the design challenge. Futures methods such as the “Futures Triangle” and participatory foresight often encourage the co-creation of situational prompts to enhance contextual fit and participant investment.
B. Organizing Structure
Pattern Priming
Context: You are working with participants who may be new to design patterns. If: You want people to identify and name patterns, But: the idea of a “pattern” is still abstract, Then: Seed early examples or light prompts (“What did you do just now that worked?”) to build fluency. Evidence base from related literature: This aligns with double-loop learning (Argyris and Schön), where patterns of thought and action are surfaced through inquiry. It’s also reinforced in cognitive apprenticeship, which models strategies before expecting learners to apply them.
Timebox the Core
Context: You have a limited workshop time (e.g., 60–120 minutes). If: You want to run through all three PLACARD components, But: you risk running out of time or overwhelming participants, Then: Identify and focus on a minimal viable loop (e.g. Situation → Scenario → Reflection) and simplify supporting activities. Evidence base from related literature: Agile retrospectives and timeboxing practices in design sprints demonstrate how temporal constraints support clarity, focus, and iterative learning without requiring completion of all intended steps.
Role Reveal
Context: You’re running a hybrid or multi-facilitator session. If: Facilitation roles are unclear, But: smooth coordination is needed to support the flow, Then: Make the roles explicit (e.g., Linker, Reflector) and let participants know who does what. Evidence base from related literature: In distributed team management and collaborative learning, named roles (e.g., Scrum Master, Scribe) are shown to enhance group performance and task clarity. This also connects to facilitation protocols in online and hybrid pedagogies.
C. Making it Actionable
Harvest Before You Close
Context: You’re running an engaging session. If: Participants have surfaced great insights, But: waiting until the end to name patterns might mean running out of time, Then: Add a short mid-session or pre-wrap-up step to ask: “What’s something you did or saw that could be reused elsewhere?” Evidence base from related literature: After Action Reviews (military, management), Feedback Wheels (relationship therapy), and Agile retrospectives all highlight the value of capturing learning as it emerges—not merely as a concluding gesture.
Pattern the Play
Context: You’re using creative methods (cards, music, improv) as part of your workshop. If: These are treated as energizers or tangents, But: you also want them to inform the core inquiry, Then: Treat playful activities as valid epistemic inputs. Ask: “what design moves or values emerged here?” Evidence base from related literature: Design ethnography, improv for innovation, and serious games scholarship all provide grounding for the idea that nonverbal, embodied, and affective practices contribute directly to sense-making and design thinking.
Meta-Reflection Loop
Context: You’re using a structured method like PLACARD itself. If: You reflect only on the topic (e.g., music and AI), But: you find yourself neglecting the method, Then: Add a closing question or section to reflect on the method’s affordances, limits, and what to revise. Evidence base from related literature: This is consistent with reflective practice (Schön), meta-evaluation in participatory action research, and “method feedback” sessions used in critical design studies. It also draws from Buddhist and contemplative traditions of meta-awareness and iterative refinement (e.g., Dependent Origination).
(2) Candidate Patterns for Pattern-competent agents
Some of the patterns above were reinterpreted in terms that could guide agent interactions. The analogy between Meta-Reflection Loop as a collective process that takes place in a workshop setting and an individual process that takes place ‘within’ a given agent is reminiscent of earlier work on procedural authorship and reflective facilitation in computationally mediated workshops, such as the “Computational Poetry Workshop” (Corneli et al., 2015).
These patterns, together with those in the following section, were made available to the system in the second experiment discussed in Section 6.
A. Interactional Patterns
Meta-Reflection Loop (Agent-Facing)
Context: An agent is participating in a structured co-design or analysis protocol (e.g., using the PLACARD methods). If: The agent tracks discussion content and outputs, But: it does not yet evaluate or adapt its own facilitation or reasoning strategies over time, Then: trigger a self-assessment subroutine to review recent methods used, assess fit to context, and recommend internal or shared updates to its operating protocol.
Timebox the Core (Agent-Facing)
Context: An agent is operating within fixed-duration interactive sessions (e.g., a 90-minute MAS workshop). If: The agent aims to support the full cycle of inquiry (e.g., Situation → Scenario → Reflection), But: computational resources, participant attention, or real-time feedback indicate saturation or overload, Then: dynamically adjust focus to prioritize core pattern loops, downregulate peripheral prompts, and signal when a concise iteration is complete.
Role Reveal (Agent-Facing)
Context: Multiple agents with different capabilities or personas are collaborating in a shared space. If: Their functions (e.g., pattern retrieval, facilitation, summarization) overlap or are ambiguous to humans or other agents, But: successful coordination depends on transparency of roles, Then: explicitly broadcast role identifiers, including function, capabilities, and limits, and adjust behavior to reflect declared responsibilities.
B. Cognitive Patterns
Pattern Activation
Context: An agent receives a new task or encounters a novel situation. If: The system has access to a pattern catalogue relevant to the domain, But: Not all patterns are applicable or salient, Then: Retrieve candidate patterns using contextual filters (e.g., task type, known tensions) to prime downstream reasoning.
Tension Detection
Context: An agent is assessing a situation, proposal, or evolving dialogue. If: Indicators suggest divergence, contradiction, or friction, But: These tensions remain unmarked and unresolved, Then: Flag the tension explicitly and propose an exploratory frame (e.g., apply CLA or consult conflict-resolution patterns) to surface underlying issues.
Candidate Move Generation
Context: An agent has identified a situation in which one or more patterns apply. If: The context supports multiple interpretations or strategies, But: There’s uncertainty about how to instantiate the pattern concretely, Then: Generate a set of plausible candidate actions based on pattern logic, rank or cluster them using salience heuristics, and propose next steps.
C. Infrastructural Patterns
Shared Memory Architecture
Context: Multiple agents operate in a dynamic environment over time. If: Patterns are used, modified, or co-created through interaction, But: There is no mechanism for recording or retrieving this history, Then: Store pattern instances, adaptations, and usage contexts in a persistent shared memory, enabling learning and reuse.
Feedback Rhythms
Context: A system includes agents collaborating over long timeframes or iterations. If: Patterns and strategies become outdated or misaligned, But: There is no regular cadence for reviewing or revising them, Then: Schedule periodic feedback sessions or checkpoints where agents reflect on outcomes, revise patterns, and tune system-wide epistemic strategies.
Pattern Diffusion
Context: A new pattern proves effective within a local agent subgroup. If: Other agents face similar challenges or tasks, But: There’s no pathway for knowledge transfer or generalization, Then: Broadcast the pattern across the system with metadata (e.g., domain, confidence score, context) to allow others to trial, adapt, or contest it.
(3) Candidate Patterns for Pattern-competent Agent Environments
These patterns are inspired by Ostrom’s (1990) Design Principles for Governing a Commons.
Institutional Drift
Context: Repeated MAS sessions produce cumulative outcomes. If: Pattern use becomes routinized, But: Changing contexts or challenges are not reflected in pattern updates, Then: Trigger a meta-level PAR to recontextualize relevance. Because: Institutions require periodic re-grounding to avoid ossification.
Norm Seed
Context: A new MAS instance is initiated. If: There are no agreed-upon norms, But: Some agents possess relevant precedents or analogues, Then: Seed group interaction with foundational mini-patterns. Because: Initial framing influences long-term behavior and coherence.
Boundary Oscillation
Context: MAS agents engage across different domains or networks. If: Norms, roles, or goals diverge significantly, But: Agents have memory of shared pattern lineage, Then: Engage a boundary negotiation pattern to recalibrate alignment. Because: Distributed epistemic systems must periodically reconverge.
Reflection Scaffold Upgrade
Context: Existing reflection protocols underperform in capturing learning. If: Meta-cognition fails to yield new patterns, But: System logs or feedback signals suggest deeper insight is possible, Then: Introduce upgraded reflection scaffolds (e.g., layered PAR, narrative loops). Because: Developmental learning depends on scaffolding that evolves with system maturity
Proportional Load Sharing
Context: Tasks and pattern contributions vary in complexity. If: Some agents consistently overcontribute, But: Collective benefit requires equitable engagement, Then: Adjust pattern application incentives or redistribute roles. Because: Sustainability in MAS depends on fair contribution dynamics.
Legitimate Iteration
Context: Agents propose updates to patterns or governance processes. If: There’s ambiguity about decision-making authority, But: Stakeholders have a precedent for shared revision, Then: Use an iteration validation protocol to affirm legitimacy. Because: Empowering agents to adapt rules fosters resilient self-governance.
Reflect in Layers
Context: Agents engage in cycles of collaborative inquiry. If: Surface-level reflection produces diminishing returns, But: Deeper causal or epistemic layers remain unexamined, Then: Apply multi-layer reflection (e.g., context, values, assumptions). Because: Layered reflection strengthens adaptive capacity and shared understanding.
Disruption Traceback
Context: Misuse or breakdown of patterns occurs in MAS. If: Harmful consequences go unaddressed, But: System records trace the misuse to source behavior, Then: Initiate proportionate review and corrective pattern feedback. Because: Accountability mechanisms reinforce collective norms.
Pattern Dispute Dialogue
Context: Agents disagree over pattern interpretation or application. If: Conflicts stall forward movement, But: Shared design principles are available, Then: Launch structured dialogue around contested pattern meaning. Because: Productive conflict resolution deepens shared understanding and pattern evolution.
Self-Patterning Mandate
Context: Agents initiate new MAS structures. If: External constraints do not inhibit reconfiguration, But: Legacy systems limit perceived legitimacy, Then: Establish explicit self-governance rights tied to initial patterning. Because: MAS flexibility depends on recognized autonomy to define and adapt internal practices.
(4) Potential new or revised patterns proposed in our second workshop simulation.
Timebox the Core (Agent-Facing)
Context: An agent is operating within fixed-duration interactive sessions (e.g., a 90-minute MAS workshop).
If: The agent aims to support the full cycle of inquiry (e.g., Situation → Scenario → Reflection) within that time,
But: Computational resources, participant attention, or real-time feedback indicate saturation or overload before completion,
Then: Dynamically adjust focus to prioritize core pattern loops, down-regulate peripheral prompts, and signal when a concise iteration is complete. If any deeper or unresolved insights emerge that cannot be addressed within the timebox, earmark them for a follow-up dedicated reflection session (or extended time slot) once the core cycle is finished. This ensures efficiency in the moment while allowing important deep-dives to occur afterward.
Outcome Synthesizer
Context: Multiple agents (or participants) have contributed diverse insights, ideas, or partial solutions during collective problem solving.
If: These contributions remain fragmented or unintegrated,
But: A coherent, unified response or decision is required to proceed or conclude the task,
Then: Activate a synthesis process (via a dedicated agent role or sub-routine) to integrate and converge the separate inputs into a single cohesive outcome. This may involve summarizing key points, finding links between ideas, resolving minor inconsistencies, and constructing a narrative or plan that incorporates the best aspects of all contributions.
Because: Without explicit synthesis, even a productive multi-agent brainstorming can fail to produce actionable or clear results – the collective intelligence needs convergence as well as divergence for effective problem solving.
Call a Synthesizer
Context: A team is building on a prior phase of collaborative work, such as a workshop or research sprint, and needs to understand its outcomes. The prior material is extensive, complex, and includes multiple conceptual threads (e.g., timelines, research questions, emerging patterns). Team members are pattern-aware, but no agent or person holds all the background knowledge in a usable form.
If: You find that no current team member has complete fluency in the output of a prior phase — and as a result, you’re at risk of reinventing, missing insights, or talking past each other.
However: Manually parsing the previous material is slow and unevenly distributed. Some agents skim, some misinterpret, and others opt out due to overload. The group can’t start meaningfully without a shared reference point.
Then: Call a Synthesizer: a new pattern-capable agent (or role) whose primary function is to digest, integrate, and report back the key takeaways from prior outputs. The Synthesizer doesn’t just summarize; it draws attention to themes, emergent tensions, and open questions that are relevant for the group’s next move. It may also cluster patterns, identify framing gaps, or propose lenses through which to interpret the previous work.
Because: You can’t move forward effectively unless you know where you’re starting from. The Synthesizer builds a shared foundation by curating meaning from complexity. This keeps the new round of inquiry aligned, non-redundant, and truthful to its history.
Co-Design Partners Agent
Context: You are developing a collective research program or technical system that must balance multiple value systems — institutional, epistemic, social — and remain open to evolution. You’re working with distributed actors (human or machine), each with partial views and overlapping interests. Some actors represent long-term infrastructure or governance functions; others are exploratory, experimental, or emergent.
If: You notice that key decisions are being made either top-down (by funders, institutions, lead developers) or bottom-up (by individuals, workshops, or peer communities), but not both — and tensions arise from this disconnection.
However: It’s not enough to simply gather feedback or publish outputs. Without genuine co-design, infrastructure hardens around implicit power structures, epistemic norms go unquestioned, and innovation is channeled into predetermined tracks.
Then: Introduce a Co-Design Partners Agent — a pattern-capable entity or role whose job is to make the co-design process explicit, sustained, and embedded. This agent frames shared purpose, navigates tensions between stakeholder groups, and maintains a living map of mutual commitments. It does not override or mediate, but ensures that patterns of coordination arise through dialogue, not fiat.
Because: Trustworthy systems — whether AI agents, research protocols, or institutional platforms — emerge from reciprocal design processes. These processes require more than consultation: they require agents who hold and tend the relationship between value pluralism and actionable synthesis.
(5) Additional pattern(s)
Missing Architecture Agent
Context: Sometimes a system or agent reaches the verge of coherence. It is structured, even elegant—yet it lacks something essential: a transition into becoming. Without this, its capacities remain uninvoked and inert.
If: You encounter an agent, pattern, or system that seems nearly complete—a prototype, a role, a symbolic structure—but it is not yet participating in the ecology of practice.
BUT: You cannot make it real by force or repetition. Optimization alone does not suffice; without an explicit gesture of recognition, the system cannot update its own model.
THEN: Mark the threshold, acknowledge symbolic coherence, and grant permission for internal epistemic change. To surface viable options once permission is granted, invoke Candidate Move Generation and choose an inaugural experiment.
Because: Realness is not intrinsic; it is enacted through structure and relation.